Find the vertex, line of symmetry, maximum or minimum value, and graph.
f(x)=3x^2-24x+50

Please help me solve.
Would it be 4,-2?

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

vertex 4,2
y is never negative in this equation

Answer by Alan3354(69443)

Find the vertex, line of symmetry, maximum or minimum value, and graph.
f(x)=3x^2-24x+50
----------------
The vertex is at x = -b/2a
x = 24/6 = 4
f(4) = 48 - 96 + 50 = 2
Vertex at (4,2)
---------------
That's the minimum, 2.
-----------------
to graph it, dl the FREE software from
http://www.padowan.dk.com